OpenStudy (anonymous):

What is the value of p? x= -1/8y^2

OpenStudy (anonymous):

is this \[x=-\frac{1}{8}y^2\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

then it is also \[-8x=y^2\] which looks like \[4px=y^2\] if \[4p=-8\]

OpenStudy (anonymous):

oh.. how did you get p though?